# Quantitative surgery and total mean curvature

### v1

@@ Abstract.
We develop quantitative surgery, which extends the classical constructions of Gromov–Lawson and Lawson–Michelsohn. As an application, we prove a conjecture of Gromov on the total mean curvature of fill-ins.
@@ 1. Introduction
Let $(M,g_{M})$ be a closed Riemannian manifold, possibly disconnected, with scalar curvature $\mathrm{scal}_{g_{M}}\colon M\to\mathbb{R}$. A fill-in of $(M,g_{M})$ is a compact Riemannian manifold $(\Omega,g_{\Omega})$ such that $(\partial\Omega,g_{\partial\Omega})$ is isometric to $(M,g_{M})$, where $g_{\partial\Omega}$ denotes the metric induced by $g_{\Omega}$ on $\partial\Omega$. We write $H_{g_{\Omega}}\colon\partial\Omega\to\mathbb{R}$ for the mean curvature of $\partial\Omega\subset(\Omega,g_{\Omega})$, where we adopt the convention that the mean curvature of the closed unit ball in flat Euclidean space $\mathbb{R}^{n}$ is equal to n1n-1.
Gromov formulated the following conjecture:
Conjecture 1.1 ([18, p. 232]).
@@
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma).$
The following is our main result.
Theorem 1.2.
@@
Let $(M,g_{M})$ be a closed Riemannian manifold admitting a spin structure. Then, for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying:
  • $\mathrm{scal}_{g_{\Omega}}\geq\sigma$,
  • $H_{g_{\Omega}}\geq 0$,
  • $\Omega$ admits a spin structure in case $\dim M\geq 7$,
one has:
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma).$
@@
Theorem 1.2 has been proven before in the following cases: By Shi–Wang–Wei [25] if $M$ is diffeomorphic to the $n$-sphere, based upon Shi–Tam’s proof for the positivity of the Brown–York mass [24] and the hyperbolic positive mass theorem (PMT) [29, 14, 2]; by Wang [30] for tori $T^{n}$, using Brendle–Hung’s [11] work on the Horowitz–Myers PMT; and by Bär in [5] concurrently under the additional assumption that $\Omega$ admits a spin structure.
Our approach, which differs from the previously mentioned works, relies on performing surgeries of sufficiently high codimension to reduce Theorem 1.2 to the case where $M$ is a sphere, which is covered in [25]. In particular, no spin assumption on the fill-in manifold $\Omega$ is required provided the hyperbolic positive mass theorem holds for orientable manifolds in dimension $\dim M+1$.
@@
The spin assumption on $M$ in Theorem 1.2 is used solely to guarantee the existence of the required surgeries transforming $M$ into a sphere, whereas the spin assumption on $\Omega$ is only needed for the validity of the hyperbolic positive mass theorem. In particular, our argument does not rely on index theory for a spin Dirac operator on either $M$ or $\Omega$. In fact, if 1.1 is proven for one simply connected, totally non-spin manifold in dimension $n$, e.g. $\mathbb{CP}^{2}\times S^{n-4}$ or the non-trivial linear $S^{n-2}$-bundle over $S^{2}$, then our proof of Theorem 1.2 shows that 1.1 holds true for every orientable manifold $M$ of dimension $n$.
@@ 1.1. Quantitative surgery
Our main tool for establishing Theorem 1.2 is a new quantitative surgery procedure. This encompasses and strengthens the classical results of Gromov–Lawson [17] and Lawson–Michelsohn [20].
To explain this construction, let $S$ be a compact submanifold in $(M,g_{M})$ of codimension $k$, and let $\varrho>0$ be a small parameter. A Gromov–Lawson handle of size $\varrho$ is a hypersurface $\Sigma_{\varrho}$ contained in $(M\times[0,1],dt^{2}+g_{M})$ which coincides with $M=M\times\{0\}$ outside the $2\sqrt{\varrho}$-tubular neighbourhood of $S$ in $M$, becomes cylindrical of width $\varrho^{4}$ at height $2\varrho$ and is invariant under the fibrewise $O(k)$-symmetry of the normal bundle of $S$ in $M$, see Fig. 1.
Let $\mathrm{scal}_{\varrho}$ be the scalar curvature and let $H_{\varrho}$ be the mean curvature of $\Sigma_{\varrho}$. Furthermore, let $s$ be the minimum of the scalar curvature of $(M,g_{M})$ on $S$. The crucial properties of our construction are:
  • $(H_{\varrho})_{-}$ converges to zero as $\varrho\to 0$.
  • $(\mathrm{scal}_{\varrho}-s)_{-}$ converges to zero as $\varrho\to 0$.
Consequently, by performing a conformal transformation arbitrarily close to the identity, we can ensure that the mean curvature is positive for sufficiently small $\varrho$.
The detailed construction in the special case where $S$ is a point in flat Euclidean space $\mathbb{R}^{n}$ is carried out in Section 4. For general $M$ and $S$, error terms arise, which will be controlled in Section 5.
@@
Figure 1. A cross-section of a family of Gromov–Lawson handles $\Sigma_{\varrho}$ attached to $\Omega$ along $M=\partial\Omega$. The origin corresponds to the intersection of the cross-section with $S$.
@@
Our construction moreover provides precise control between handles of different sizes $\varrho$. In particular, the heuristic observation (see Figure 1) that the metrics on $\Sigma_{\varrho}$ are monotonically increasing in $\varrho$ can be made rigorous, while maintaining a uniform lower bound on the scalar curvature of $\Sigma_{\varrho}$. This feature is crucially exploited in the proof of Theorem 1.2, as described below. Our computations are fully explicit, and the handles are constructed from first principles, yielding an independent and unifying proof of both [17, Theorem A] and [20, Theorem 3.1].
@@
Owing to these explicit geometric bounds, our results also encompass variants of the conclusions of several earlier works, including [26, 27, 7, 15, 9, 8].
1.2. Proof outline
By topological arguments, we can assume that we can pass from $M$ to the sphere $S^{n}$ through surgeries of codimension $k\geq 3$. At the same time, we pass from a fill-in of $M$ to a fill-in of $S^{n}$ via the corresponding handle attachments. Working with handles of sufficiently small width, this ensures that the total mean curvature of the original fill-in of $M$ transfers to the total mean curvature of a fill-in of $S^{n}$, up to a small error. An application of Shi–Wang–Wei’s total mean curvature estimate to $S^{n}$ then concludes the proof of Theorem 1.2.
Relying on the classical surgery result of Lawson–Michelsohn [20], this strategy can be carried out as stated when restricting to fill-ins $\Omega$ of $M$ with $H_{g_{\Omega}}\geq\mu$ for some fixed $\mu>0$ and working with a constant $\Lambda$ that depends on $M$, $g_{M}$, $\sigma$ and $\mu$. We present the proof of this special case of Theorem 1.2 in Section 2.
In the general case, however, there is a significant caveat. If the point-wise lower mean curvature bound of different fill-ins $\Omega$ is not uniformly bounded below by a positive constant, the handles have to be chosen thinner and thinner to ensure a non-negative mean curvature bound for the resulting fill-in of $S^{n}$. This leads to infinitely many metrics on the resulting $S^{n}$ and consequently, Shi–Wang–Wei’s result does not yield a uniform total mean curvature bound. At this stage, it is therefore unclear whether the constant $\Lambda$ can be chosen independently of $\mu$.
This difficulty will be overcome by the following strategy (see Section 3): Shi–Wang–Wei’s upper bound can be chosen uniformly for a family of metrics on $S^{n}$, provided that each member of the family can be connected to a fixed metric on $S^{n}$ along a monotonically increasing path of Riemannian metrics with lower scalar curvature bounds that are independent of the given metric in the family (see Proposition 3.2 for the precise statement). Our quantitative surgery construction in Section 1.1 provides the required control of the surgery handles. Proceeding by induction on the number of surgery steps finishes the proof in the general case.
@@ 1.2. Proof outline
Theorem 1.2 also allows us to define a new quasi-local mass with several particularly appealing properties. This will be discussed in Section 7.
@@
Acknowledgements: BH and SH are grateful to the University of Augsburg, to Columbia University, the Lonavala Geometry Festival, and to the Chennai Mathematical Institute for their hospitality and ideal working conditions. BH acknowledges support by the DFG-SPP 2026 “Geometry at Infinity”. SH was supported in part by an AMS–Simons Travel Grant (award no. 210022).
@@
2. Total mean curvature bounds with fixed positive lower mean curvature bound
Here we will prove the following weak version of Theorem 1.2.
Theorem 2.1.
Let $(M,g_{M})$ be a closed Riemannian manifold admitting a spin structure. Let $\mu>0$.
Then, for every $\sigma\in\mathbb{R}$, there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma,\mu)\in\mathbb{R}_{+}$ such that for every fill-in $(\Omega,g_{\Omega})$ of $(M,g_{M})$ satisfying:
  • $\mathrm{scal}_{g_{\Omega}}\geq\sigma$,
  • $H_{g_{\Omega}}\geq\mu$,
  • if $\dim M\geq 7$, then $\Omega$ is a spin manifold and there is a spin structure preserving isometry $(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$,
one has:
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{M}}\leq\Lambda(M,g_{M},\sigma,\mu).$
In the remainder of this section, we prove this result. Put $n:=\dim M$. Let $\sigma\in\mathbb{R}$ and let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ satisfying the properties of Theorem 2.1. For simplicity of exposition, we assume $(M,g_{M})=\partial(\Omega,g_{\Omega})$. We start with some preliminary topological modifications. By taking products of $M$ and of $\Omega$ with standard round spheres, we can assume without loss of generality that $n\geq 5$. Upon passing to the oriented double cover of $(\Omega,g_{\Omega})$ and to the trivial double cover of $(M,g_{M})$, we can furthermore assume that $\Omega$ is orientable.
The kernel of the forgetful map from spin bordism to oriented bordism is a $2$-torsion group by [1, Corollary 2.6]. Replacing $M$ by $M\sqcup M$ and $\Omega$ by $\Omega\sqcup\Omega$, we can hence assume that $M$ bounds a connected spin manifold $B$. Furthermore, if $\dim M\geq 7$, then the given spin structure on $\Omega$ extends over $\Omega\cup_{M}B$.
@@ 2. Total mean curvature bounds with fixed positive lower mean curvature bound
Removing a small open disc from the interior of $B$, we obtain a compact connected spin bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{1}(W)$ and $\pi_{2}(W)$ by surgeries in the interior of $W$ (and using $n\geq 5$), we can assume that $W$ is $2$-connected. In particular, the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $2$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $3$.
@@
Let $\ell\geq 0$ be the number of handles in this handle decomposition of $W$. Using this handle decomposition, $S^{n}$ is obtained from $M$ by $\ell$ spin surgeries of codimension at least $3$. The proof of Theorem 2.1 is by induction on $\ell$. The case $\ell=0$ is covered by [25, Theorem 4.1 and Remark 4.2]. Now assume by induction that Theorem 2.1 holds for some $\ell\geq 0$ and assume that the number of codimension $3$-handles in $W$ is equal to $\ell+1$.
@@
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$ be the surgery datum for the first codimension $3$-handle in $W$ and let $\widetilde{M}$ be obtained from $M$ by surgery along $\psi$. Note that the induction hypothesis applies to $\widetilde{M}$. There exist smooth embeddings $\psi_{i}\colon S^{n-k}\times D^{k}\hookrightarrow M$, $i=1,2$, that are both isotopic to $\psi$ and have disjoint images, see Fig. 2. The point of this “doubling construction” is to estimate the total mean curvature after the surgery by a pigeonhole argument, see (2.2) below.
@@
Choose a smooth collar $M\times[0,1)\hookrightarrow W$ of $M\subset\partial W$ in $W$. With respect to this collar, we identify $M\times[0,1)$ with a smooth submanifold of $W$. Choose a Riemannian metric $h$ on $W$ which restricts to the metric
(2.1) $dt^{2}+\left(1+\tfrac{t}{n-1}\mu\right)g_{M}\text{ on }M\times[0,1)\subset W,$
where $g_{M}$ and $\mu$ are from the assumptions in Theorem 2.1. Note that $(W,h)$ only depends on $(M,g_{M})$, but not on the fill-in $(\Omega,g_{\Omega})$. The second fundamental form of $(W,h)$ along $M\times\{0\}$ with respect to the exterior normal is equal to $\tfrac{\mu}{2(n-1)}\cdot g_{M}$.
From now on, let $i\in\{1,2\}$. By [20, Theorem 3.1], we can find a submanifold $T_{i}\subset W$ which intersects $M$ in $\mathrm{im}(\psi_{i})$, which is isotopic to the handle attached along the surgery datum $\psi_{i}$, such that $\Omega\cup T_{i}$ is smooth and such that the mean curvature
$\widetilde{M}_{i}:=\partial\left(\Omega\cup T_{i}\right)$
is bounded below by some constant $0<\mu_{i}<\tfrac{\mu}{2}$, see Fig. 3. Here $\partial\left(\Omega\cup T_{i}\right)$ is regarded as a smooth submanifold of $W$ and the mean curvature is computed with respect to the metric $h$ and the normal pointing inside $W$. Note that $\widetilde{M}_{i}$ carries a spin structure induced by the one on $W$.
@@
Figure 3. The cobordism $W$ consisting of one handle and the embedded handle
Let $\widetilde{g}_{i}$ be the metric on $\widetilde{M}_{i}$ induced by $h$. Note that the induction hypothesis applies to $(\widetilde{M}_{i},\widetilde{g}_{i})$, as $\widetilde{M}_{i}$ is diffeomorphic to $\widetilde{M}$.
Let $U_{i},V_{i},Z_{i}$, $i=1,2$, be open neighborhoods of $\mathrm{im}(\psi_{i})\subset M$ such that $\overline{U}_{i}\subset V_{i}$, $\overline{V}_{i}\subset Z_{i}$ and $\overline{Z}_{1}\cap\overline{Z}_{2}=\emptyset$.
By [4, Theorem 3.7], we can smoothly deform the metric $g_{\Omega}$ on $\Omega$ to a new metric on $\Omega$, called $g_{i}$ for short, with the following properties (see Fig. 4):
  • The metrics on $\partial\Omega$ induced by $g_{i}$ are equal to $g_{M}$,

### v2

# Surgery and total mean curvature

@@ Abstract.
We prove Gromov’s conjecture on the total mean curvature of fill-ins in various cases. Our methods are based on surgery to reduce the statement to fill-ins of spheres, which can be treated by instances of the positive mass theorem. For spin fill-ins, where we permit the mean curvature to take negative values, we build on a classical surgery result of Lawson–Michelsohn and a recent positive mass theorem with creases by Kazaras–Khuri–Lin. For non-spin fill-ins of spin manifolds, where we assume the mean curvature to be non-negative, we develop a novel quantitative surgery process to reduce the general situation to a result of Shi–Wang–Wei. We also treat the case of fill-ins of non-spin manifolds, provided there is a fixed positive lower bound on the mean curvature.
Key words and phrases:
Fill-ins, scalar curvature, total mean curvature, surgery; positive mass theorem
@@ 1. Introduction
Let $(M,g_{M})$ be a closed Riemannian manifold, possibly disconnected, with scalar curvature $\mathrm{scal}_{g_{M}}\colon M\to\mathbb{R}$. A fill-in of $(M,g_{M})$ is a compact Riemannian manifold $(\Omega,g_{\Omega})$ such that $(\partial\Omega,g_{\partial\Omega})$ is isometric to $(M,g_{M})$, where $g_{\partial\Omega}$ denotes the metric induced by $g_{\Omega}$ on $\partial\Omega$. We write $H_{g_{\Omega}}\colon\partial\Omega\to\mathbb{R}$ for the mean curvature of $\partial\Omega\subset(\Omega,g_{\Omega})$ with respect to the exterior normal. By convention, this means that the mean curvature of the closed unit ball in flat Euclidean space $\mathbb{R}^{n}$ is equal to n1n-1.
Inspired by the positive mass theorem in general relativity, Gromov conjectured that for fixed $(M,g_{M})$, the total mean curvature of $(\Omega,g_{\Omega})$ cannot be arbitrarily large under a uniform lower scalar curvature bound of $(\Omega,g_{\Omega})$:
Conjecture 1.1 ([Gro23, p. 232]).
@@
$\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq\Lambda(M,g_{M},\sigma).$
The following result verifies this conjecture for various cases, with a constant that also depends on a lower mean curvature bound.
Theorem.
@@
Let $(M,g_{M})$ be a closed Riemannian manifold of dimension $n$ and let $\sigma,\kappa\in\mathbb{R}$. Then there exists a constant $\Lambda=\Lambda(M,g_{M},\sigma,\kappa)>0$ with the following property: If $(\Omega,g_{\Omega})$ is a fill-in of $(M,g_{M})$ such that:
  • $H_{g_{\Omega}}\geq\kappa$,
  • $\mathrm{scal}(g_{\Omega})\geq\sigma$,
and such that one of the following conditions holds:
  1. (A)
    $\kappa>0$ and $n\leq 6$,
  2. (B)
    $\kappa=0$, $n\leq 6$ and $M$ admits a spin structure,
  3. (C)
    $\Omega$ admits a spin structure,
then
$\int_{M}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq\Lambda(M,g_{M},\sigma,\kappa).$
The dependence of $\Lambda$ on the Riemannian metric $g_{M}$ is unavoidable, even if $\mathrm{vol}(M)$ and $\mathrm{diam}(M)$ are uniformly bounded, see Fig. 1 and Appendix B. The explicit dependence of $\Lambda$ in terms of $\sigma$ and $\kappa$ is worked out in (3.8).
Figure 1. A Riemannian manifold with large $\Lambda(M,g_{M},0,0)$.
@@
Assuming $\kappa\geq 0$, this result has been proven before in the following cases: By Shi–Wang–Wei [SWW22] if $M$ is diffeomorphic to the $n$-sphere and for orientable fill-ins, based upon Shi–Tam’s proof for the positivity of the Brown–York mass [ST02] and the hyperboloidal positive mass theorem (PMT) [Wan01, CJŁ04, ACG08]; and by Wang [Wan24] for tori $T^{n}$ and fill-ins $D^{2}\times S^{n-1}$, using Brendle–Hung’s [BH24] work on the Horowitz–Myers PMT.
Christian Bär in [Bär26] has presented a different proof of Theorem (C) using spinorial methods. Theorem (C) and Bär’s result provide the first upper bounds on the total mean curvature to treat fill-ins whose mean curvature assumes negative values.
Convention.
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ as in Theorem (A), Theorem (B) or Theorem (C). Then the orientation double cover $(\overline{\Omega},g_{\overline{\Omega}})$ of $(\Omega,g_{\Omega})$ is a fill-in of the orientation double cover of $(M,g_{M})$ and the assumptions in the respective theorem are still satisfied. Furthermore, $\int_{\partial\overline{\Omega}}H_{g_{\overline{\Omega}}}\mathrm{dvol}_{g_{\partial\overline{\Omega}}}=2\int_{\partial\Omega}H_{g_{\Omega}}\mathrm{dvol}_{g_{\partial\Omega}}$. Therefore, in our proofs, we can restrict to oriented $(M,g_{M})$ and $(\Omega,g_{\Omega})$ with an orientation preserving Riemannian isometry $(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$. These assumptions will be used throughout the rest of the paper.
1.1. Proof outline
Our approach is different from the previous ones. By topological arguments, we can assume that we can pass from $M$ to the sphere $S^{n}$ through surgeries of codimension $k\geq 2$. At the same time, we pass from a fill-in of $M$ to a fill-in of $S^{n}$ via the corresponding handle attachments. Working with handles of sufficiently small width and a pigeonhole argument, this ensures that the total mean curvature of the original fill-in of $M$ transfers to the total mean curvature of a fill-in of $S^{n}$, up to a controlled error.
Relying on the classical surgery result of Lawson–Michelsohn [LM84] for positive mean curvature, this strategy can be carried out as stated when restricting to fill-ins $\Omega$ of $M$ with $H_{g_{\Omega}}\geq\kappa$ for some fixed $\kappa>0$. Using the result of Shi–Wang–Wei [SWW22], this leads to a proof of Theorem (A), which we present in Section 2.
The same strategy also applies to Theorem (C), when working with a positive mass theorem for manifolds with creases due to [Lin23, KKL25]. The surgery part is even simpler in this case, as the result of Lawson–Michelsohn is not needed.
From a technical point of view, the proof of Theorem (B) is the most challenging task. The majority of this paper is devoted to addressing this issue. Using surgery, we again reduce the general case to $M=S^{n}$ treated by Shi–Wang–Wei. However, if the pointwise lower mean curvature bound of different fill-ins $\Omega$ is not uniformly bounded below by a positive constant, the handles have to be chosen thinner and thinner to ensure that the mean curvature for the resulting fill-in of $S^{n}$ is nonnegative, which is required in [SWW22]. This leads to infinitely many metrics on the resulting $S^{n}$ and consequently, Shi–Wang–Wei’s result does not yield a uniform total mean curvature upper bound.
This difficulty is overcome by the following strategy (see Section 4): Shi–Wang–Wei’s upper bound can be chosen uniformly for a family of metrics on $S^{n}$, provided that each member of the family can be connected to a fixed metric on $S^{n}$ along a monotonically increasing path of Riemannian metrics with lower scalar curvature bounds that are independent of the given metric in the family (see Proposition 4.2 for the precise statement). We will develop a quantitative surgery construction, see Section 1.2, that provides the required control of the surgery handles. Proceeding by induction on the number of surgery steps finishes the proof for Theorem (B).
@@ 1.1. Proof outline
Theorem (A) and Theorem (B) hold for any $n$ for which the hyperboloidal PMT holds for oriented manifolds in dimension $n+1$.
@@ 1.2. Quantitative surgery
Our main tool for establishing Theorem (B) is a new quantitative surgery procedure. This encompasses and strengthens the classical surgery result of Gromov–Lawson [GL80] and has connections to the extrinsic surgery result of Lawson–Michelsohn [LM84].
To explain this construction, let $\mathcal{S}$ be a compact submanifold in $(M,g_{M})$ of codimension $k$, and let $\varrho>0$ be a small parameter. A Gromov–Lawson handle of size $\varrho$ is a hypersurface $\Sigma_{\varrho}$ contained in $(M\times[0,1],dt^{2}+g_{M})$ which coincides with $M=M\times\{0\}$ outside the $2\sqrt{\varrho}$-tubular neighbourhood of $\mathcal{S}$ in $M$, becomes cylindrical of width $\varrho^{4}$ at height $2\varrho$ and is invariant under the fibrewise $\mathrm{O}(k)$-symmetry of the normal bundle of $\mathcal{S}$ in $M$, see Fig. 2.
Let $\mathrm{scal}_{\varrho}$ be the scalar curvature of $\Sigma_{\varrho}$ and let $H_{\varrho}$ be the mean curvature of $\Sigma_{\varrho}$. Furthermore, let $s$ be the minimum of the scalar curvature of $(M,g_{M})$ on $\mathcal{S}$. The crucial properties of our construction are:
  • if $k\geq 2$, then $(H_{\varrho})_{-}$ converges to zero as $\varrho\to 0$,
  • if $k\geq 3$, then $(\mathrm{scal}_{\varrho}-s)_{-}$ converges to zero as $\varrho\to 0$.
By performing a conformal transformation arbitrarily close to the identity, we can ensure that the mean curvature is positive for sufficiently small $\varrho$.
The detailed construction in the special case where $\mathcal{S}$ is a point in flat Euclidean space $\mathbb{R}^{n}$ is carried out in Section 5. For general $M$ and $\mathcal{S}$, error terms arise, which will be controlled in Section 6.
@@
Figure 2. A cross-section of a family of Gromov–Lawson handles $\Sigma_{\varrho}$ attached to $\Omega$ along $M=\partial\Omega$. The origin corresponds to the intersection of the cross-section with $\mathcal{S}$.
@@
Our construction moreover provides precise control between handles of different sizes $\varrho$. In particular, the heuristic observation (see Figure 2) that the metrics on $\Sigma_{\varrho}$ are monotonically increasing in $\varrho$ can be made rigorous, while maintaining a uniform lower bound on the scalar curvature of $\Sigma_{\varrho}$. The last property requires working with handles of codimension at least $3$, which accounts for the spin assumption for $M$ in Theorem (B).
@@
Our computations are fully explicit, and the handles are constructed from first principles, yielding an independent approach to [GL80, Theorem A] with explicit geometric bounds. Consequently, our results also encompass variants of the conclusions of several earlier works, including [Swe25, Swe25a, BDS18, Dod20, BS21, BKS20]; compare Remark 7.2.
@@
Our results also allows us to define a new quasi-local mass in general relativity with several particularly appealing properties. This will be discussed in Section 8.
@@
Acknowledgements: BH and SH wish to thank Simon Brendle for useful discussions and his interest in this work. BH and SH are grateful to the University of Augsburg, to Columbia University, to the Lonavala Geometry Festival, and to the Chennai Mathematical Institute for their hospitality and ideal working conditions. BH acknowledges financial support by Columbia University and the DFG-SPP 2026 “Geometry at Infinity”. SH was supported in part by an AMS–Simons Travel Grant (award no. 210022).
@@
2. Surgery for fill-ins with fixed lower mean curvature bounds
Applying a surgery procedure, we will reduce the general case of Theorem (A) to (oriented) fill-ins of the $n$-sphere. Since this case has been treated in [SWW22], this finishes the proof of Theorem (A). A similar argument applies to Theorem (C), where the case of spheres will be postponed to Section 3.
Proof of Theorem (A).
Let $(M,g_{M})$ and $\sigma\in\mathbb{R},\kappa>0$ be as in Theorem (A). Put $n:=\dim M$. By taking products with standard round spheres, we can assume without loss of generality that $n\geq 5$. Let $B$ be a fixed compact oriented manifold such that $\partial B=M$ and so that $B$ induces the reverse orientation on $M$.
@@ 2. Surgery for fill-ins with fixed lower mean curvature bounds
Removing a small open disc from the interior of $B$, we obtain a compact oriented bordism $W$ from $M$ to $S^{n}$. Annihilating $\pi_{0}(W)$ and $\pi_{1}(W)$ by surgeries in the interior of $W$ (using $n\geq 5$), we can assume that the inclusion $S^{n}\hookrightarrow\partial W\subset W$ is $1$-connected. Since $n\geq 5$, this implies that $(W,M)$ has a relative handle decomposition with handles of codimension at least $2$.
@@
Let $\ell\geq 0$ be the number of handles in this handle decomposition of $W$. The proof of Theorem (A) is by induction on $\ell$. The case $\ell=0$ is covered by [SWW22, Theorem 4.1 and Remark 4.2]. Now assume by induction that Theorem (A) holds for $\ell\geq 0$ and that the number of handles in $W$ is equal to $\ell+1$.
Let $(\Omega,g_{\Omega})$ be a fill-in of $(M,g_{M})$ satisfying $\mathrm{scal}_{g_{\Omega}}\geq\sigma$ and $H_{g_{\Omega}}\geq\kappa$. Let $\omega\colon(\partial\Omega,g_{\partial\Omega})\cong(M,g_{M})$ be an (orientation preserving) Riemannian isometry. For simplicity of exposition, we will assume that $\partial\Omega=M$ and $\omega=\operatorname{\mathrm{id}}$. In the general case, the isometry $\omega$ has to be inserted when gluing subsets of $B$ to $\Omega$. This implies that $\Omega\cup_{M}B$ is oriented.
@@
Let $\psi\colon S^{n-k}\times D^{k}\hookrightarrow M$, $k\geq 2$, be the surgery datum for the first handle $T$ in $W$ and let $\widetilde{M}=\partial(\Omega\cup T)$ be obtained from $M$ by surgery along $\psi$. Note that the induction hypothesis applies to $\widetilde{M}$. There exist smooth embeddings $\psi_{i}\colon S^{n-k}\times D^{k}\hookrightarrow M$, $i=1,2$, that are both isotopic to $\psi$ and have disjoint images, see Fig. 3. With this “doubling construction” we can estimate the total mean curvature after the surgery by a pigeonhole argument, see (2.3) below.
@@
Choose a smooth collar $M\times[0,1)\hookrightarrow W$ of $M\subset\partial W$ in $W$. With respect to this collar, we identify $M\times[0,1)$ with a smooth submanifold of $W$. Let $t$ be the canonical coordinate on $[0,1)$ and choose a Riemannian metric $h$ on $W$ which restricts to the metric
(2.1) $dt^{2}+\left(1+\tfrac{2t}{n}\kappa\right)g_{M}\text{ on }M\times[0,1)\subset W,$
Note that $(W,h)$ only depends on $(M,g_{M})$, but not on the fill-in $(\Omega,g_{\Omega})$. The second fundamental form of $(W,h)$ along $M\times\{0\}$ with respect to the exterior normal is equal to $\tfrac{\kappa}{n}\cdot g_{M}$.
From now on, let $i\in\{1,2\}$. By [LM84, Theorem 3.1], we can find a submanifold $T_{i}\subset W$ which intersects $M$ in $\mathrm{im}(\psi_{i})$, which is isotopic to the handle attached along the surgery datum $\psi_{i}$ and such that the mean curvature of
(2.2) $\widetilde{M}_{i}:=\partial\left(\Omega\cup T_{i}\right)$
is bounded below by some constant $0<\widetilde{\kappa}<\kappa$, for $i=1,2$, see Fig. 4. Here $\partial\left(\Omega\cup T_{i}\right)$ is a smooth submanifold of $W$ and the mean curvature is computed with respect to the metric $h$ and the normal pointing inside $W$.
@@
Figure 4. The bordism $W$ consisting of one handle and the embedded handle
Let $g_{\widetilde{M}_{i}}$ be the metric on $\widetilde{M}_{i}$ induced by $h$. The Riemannian manifold $(\widetilde{M}_{i},g_{\widetilde{M}_{i}})$ is independent of the fill-in $(\Omega,g_{\Omega})$. Furthermore, since $\widetilde{M}_{i}$ is diffeomorphic to $\widetilde{M}$, the induction hypothesis applies to $\widetilde{M}_{i}$.
Set $\widetilde{\Omega}_{i}:=\Omega\cup T_{i}\subset\Omega\cup W$. We have $\partial\widetilde{\Omega}_{i}=\widetilde{M}_{i}$. Let $Z_{i}$ be open neighborhoods of $\mathrm{supp}(\psi_{i})\subset M$ such that $Z_{1}\cap Z_{2}=\emptyset$. Further below, we will construct Riemannian metrics $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$ such that
  1. (i)
    the metric on $\widetilde{M}_{i}$ induced by $g_{\widetilde{\Omega}_{i}}$ coincides with $g_{\widetilde{M}_{i}}$,
  2. (ii)
  3. (iii)
    $H_{g_{\widetilde{\Omega}_{i}}}\geq\widetilde{\kappa}$ everywhere,
  4. (iv)
    $\mathrm{scal}(g_{\widetilde{\Omega}_{i}})>{\rm min}\{\sigma,\min\mathrm{scal}_{h}\}-1$.
Using $Z_{1}\cap Z_{2}=\emptyset$, properties (ii), (iii) and (iv), and $\widetilde{\kappa}>0$, we obtain
(2.3) $\begin{split}\int_{\partial\Omega}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\leq{}&\int_{\partial\Omega\setminus Z_{1}}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}+\int_{\partial\Omega\setminus Z_{2}}H_{g_{\Omega}}{\rm dvol}_{g_{\partial\Omega}}\\ \leq{}&\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}}+\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\\ \leq{}&2\max\left\{\int_{\partial\widetilde{\Omega}_{1}}H_{g_{\widetilde{\Omega}_{1}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{1}}},\int_{\partial\widetilde{\Omega}_{2}}H_{g_{\widetilde{\Omega}_{2}}}{\rm dvol}_{g_{\partial\widetilde{\Omega}_{2}}}\right\}.\end{split}$
By the inductive assumption, for $i\in\{1,2\}$, Theorem (A) holds for the manifold $(\widetilde{M}_{i},\widetilde{g}_{i})$, the lower scalar curvature bound ${\rm min}\{\sigma,\mathrm{scal}_{h}\}-1$ and the lower mean curvature bound $\widetilde{\kappa}>0$. Since all of these data are independent of the fill-in $(\Omega,g_{\Omega})$, this concludes the induction step.
It remains to construct the metric $g_{\widetilde{\Omega}_{i}}$ on $\widetilde{\Omega}_{i}$. Let $U_{i},V_{i}$, $i=1,2$, be open neighborhoods of $\mathrm{im}(\psi_{i})\subset M$ such that $\overline{U}_{i}\subset V_{i}$, $\overline{V}_{i}\subset Z_{i}$. By [BH23, Theorem 3.7], we can smoothly deform the metric $g_{\Omega}$ on $\Omega$ to a new metric on $\Omega$, called $g_{i}$ for short, with the following properties (see Fig. 5):
  • The metrics on $\partial\Omega$ induced by $g_{i}$ coincide with $g_{M}$,